# The role of stacking and strain in mean-field magnetic moments of multilayer graphene

### v1

@@ II Model and methods
These lattice parameters were chosen to be consistent with the values used in previous HSE calculations [35, 36]. The two stacking arrangements differ in their interlayer registry. In ABA graphene, the second layer is shifted by the vector relative to the first layer, while the third layer is positioned directly above the first, and the fourth above the second. In ABC stacking, the first and second layers maintain the same relative positioning as in ABA, but the third layer is shifted by relative to the first layer. The unit cell structures of both systems are illustrated in the right column of Fig. 1, where even (odd) numbered sites indicate the A (B) sublattices within each layer.
@@ III Results
We now examine the magnetic properties of different multilayer graphene systems using our TB+ model with the fitted Hubbard parameter . We begin with pure stacking configurations before investigating the behaviour of mixed stacking systems. The magnetic moments of ABC multilayers calculated from our TB+ model show agreement with the hybrid DFT results, as demonstrated in Fig. 1. The magnetic structure exhibits a consistent pattern across all layer numbers: within each layer, the system displays antiferromagnetic ordering, with magnetic moment magnitudes that decay toward the center of the multilayer. Additionally, the overall magnetic moments increase systematically with the number of layers. For systems with more than 12 layers the change of the maximum value of the magnetic moments between consecutive layer number is smaller than the convergence criteria.
The ABA multilayers exhibit fundamentally different magnetic behaviour compared to their ABC counterparts. While they also show antiferromagnetic ordering within each layer (Fig. 1), the spatial distribution of magnetic moments is inverted: the magnitude increases as we move toward the center of the system. Notably, although the hybrid functional calculations predict no magnetic moments for three, four, and seven layers, our TB+ calculations reveal non-zero magnetic moments for these systems as well. Importantly, a single Hubbard parameter adequately describes the magnetic properties across both stacking configurations and multiple layer thicknesses (3-8 layers), suggesting reasonable transferability of our TB+ approach.
@@ III.1.1 Uniaxial strain in the armchair direction
We applied uniaxial strain to study its effects on the magnetic structure of pure ABC and ABA multilayers as well as selected mixed stackings. The strain was applied in the armchair () direction, parallel to one of the carbon-carbon bonds, with magnitudes of . The calculations were performed on multilayer systems with N=4, 5, 6, 9, 11 layers for both pure ABC and ABA configurations, as well as the two selected mixed stackings. The results, shown in Fig. 3, display the maximum magnetic moment values for each system to provide a clear comparison of the strain effects across different configurations. Under tensile strain (stretching), the magnitude of magnetic moments increases in ABC, ABA, and mixed stacking systems. Conversely, compressive strain exhibits asymmetric effects on the two stacking types: in ABA systems, magnetism can be suppressed by sufficiently large compression, while ABC systems maintain their magnetic character throughout the experimentally accessible strain range [52].
@@ III.1.2 Deformation of the van der Waals spacing
In addition to in-plane strain effects, we investigated how interlayer separation influences the magnetic properties by applying compression along the direction (perpendicular to the graphene planes). This allows us to probe the role of interlayer coupling in stabilizing the magnetic states. We applied changes to the interlayer distance, where corresponds to an increased layer separation of relative to the equilibrium distance Å. We chose higher strain values because compression values of up to 5% are available experimentally in van der Waals systems when hydrostatic pressure is applied [53]. The same multilayer systems (N=4, 5, 6, 9, 11 layers) were examined for both ABC and ABA stackings. The results presented in Fig. 4 show the maximum magnetic moment for each system configuration, allowing for straightforward comparison of how interlayer distance affects magnetic strength across different layer numbers and stacking types.
@@ Acknowledgements.
Supported by the DKOP-23 Doctoral Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. This work was supported by the National Research, Development and Innovation Office (NKFIH) in Hungary, through Grant No. FK-142985, K 146156 and Excellence 151372 and by the Hungarian Academy of Sciences LP2024-17 Lendület ”Momentum” grant. Z.T. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.Appendix A The tight-binding model
@@ Appendix C Simultaneous uniaxial strain and compression
In [35] and [36] the effect of different geometries was not studied. In accordance with this we fixed the lattice parameters. With these settings the TB+U model could not reproduce the absence of magnetism in 3, 4 and 7 layer ABA graphene. Calculation incorporating both uniaxial strain and interlayer compression (Fig. 9) show that for a large set of possible lattice parameters the magnetic moments are negligibly small or zero for the ABA 3 layer graphene. The same effect occurs in 5 layer ABA but for a much smaller set of lattice parameters. Meanwhile the magnetism is robust for both ABC multilayers. Therefore an explanation for the absence of magnetism in the 3,4 and 7 layer ABA systems in [36] is the selected values of the lattice parameters.

### v2

@@ II Model and methods
These lattice parameters were chosen to be consistent with the values used in previous PBE0 calculations [47, 9]. The two stacking arrangements differ in their interlayer registry. In ABA graphene, the second layer is shifted by the vector relative to the first layer, while the third layer is positioned directly above the first, and the fourth above the second. In ABC stacking, the first and second layers maintain the same relative positioning as in ABA, but the third layer is shifted by relative to the first layer. The unit cell structures of both systems are illustrated in the right column of Fig. 1, where even (odd) numbered sites indicate the A (B) sublattices within each layer.
@@ III Results
We now examine the magnetic properties of different multilayer graphene systems using our TB+ model with the fitted Hubbard parameter . We begin with pure stacking configurations before investigating the behaviour of mixed stacking systems. The magnetic moments of ABC multilayers calculated from our TB+ model show agreement with the hybrid DFT results, as demonstrated in Fig. 1. The magnetic structure exhibits a consistent pattern across all layer numbers: within each layer, the system displays antiferromagnetic ordering, with magnetic moment magnitudes that decay toward the center of the multilayer. Additionally, the overall magnetic moments increase systematically with the number of layers. The value of the magnetic moments from our TB+U calculations and [47] are reported in Table 1. For systems with more than 12 layers the change of the maximum value of the magnetic moments between consecutive layer number is smaller than the convergence criteria.
The ABA multilayers exhibit fundamentally different magnetic behaviour compared to their ABC counterparts. While they also show antiferromagnetic ordering within each layer (Fig. 1), the spatial distribution of magnetic moments is inverted: the magnitude increases as we move toward the center of the system. Notably, although the hybrid functional calculations predict no magnetic moments for three, four, and seven layers, our TB+ calculations reveal non-zero magnetic moments for these systems as well. Importantly, a single Hubbard parameter adequately describes the magnetic properties across both stacking configurations and multiple layer thicknesses (3-8 layers), suggesting reasonable transferability of our TB+ approach. The value of the magnetic moments from our work and [9] are reported in Table 2.
| N | ||||||||
| 3 | ||||||||
| 4 | ||||||||
| 5 | ||||||||
| 6 | ||||||||
| 7 | ||||||||
| 8 | ||||||||
| N | ||||||||
| 3 | ||||||||
| 4 | ||||||||
| 5 | ||||||||
| 6 | ||||||||
| 7 | ||||||||
| 8 | ||||||||
@@ III.1.1 Uniaxial strain in the armchair direction
We applied uniaxial strain to study its effects on the magnetic structure of pure ABC and ABA multilayers as well as selected mixed stackings. The strain was applied in the armchair () direction, parallel to one of the carbon-carbon bonds, with magnitudes of . Due to graphene’s physical properties’ insensitivity to the direction of strain [60] in the case of small distortions, we only study uniaxial strain in the armchair direction. The calculations were performed on multilayer systems with N=3-12 for both pure ABC and ABA configurations. The results, shown in Fig. 3, display the maximum magnetic moment values for each system to provide a clear comparison of the strain effects across different configurations. Under tensile strain (stretching), the magnitude of magnetic moments increases in ABC, ABA. Conversely, compressive strain exhibits asymmetric effects on the two stacking types: in ABA systems, magnetism can be suppressed by sufficiently large compression, while ABC systems maintain their magnetic character throughout the experimentally accessible strain range [44]. Magnetism in multilayer graphene is controlled by electronic correlation effects [17], which are determined by the ratio of in-plane to interlayer hopping amplitudes. The application of mechanical strain reduces the in-plane hopping parameters [60], thereby enhancing electronic correlations and strengthening the magnetic instability.
@@ III.1.2 Deformation of the van der Waals spacing
In addition to in-plane strain effects, we investigated how interlayer separation influences the magnetic properties by applying compression along the direction (perpendicular to the graphene planes). This allows us to probe the role of interlayer coupling in stabilizing the magnetic states. We applied changes to the interlayer distance, where corresponds to an increased layer separation of relative to the equilibrium distance Å. We chose higher strain values because compression values of up to 5% are available experimentally in van der Waals systems when hydrostatic pressure is applied [59]. The same multilayer systems N=(3-12) were examined for both ABC and ABA stackings. The results presented in Fig. 4 show the maximum magnetic moment for each system configuration, allowing for straightforward comparison of how interlayer distance affects magnetic strength across different layer numbers and stacking types.
@@ Acknowledgements.
Supported by the DKOP-23 Doctoral Excellence Program of the Ministry for Culture and Innovation from the source of the National Research, Development and Innovation Fund. This work was supported by the National Research, Development and Innovation Office (NKFIH) in Hungary, through Grant No. FK-142985, K 146156 and Excellence 151372 and by the Hungarian Academy of Sciences LP2024-17 Lendület ”Momentum” grant. Z.T. acknowledges support from the János Bolyai Research Scholarship of the Hungarian Academy of Sciences. This project is supported by the TRILMAX Horizon Europe consortium (Grant No. 101159646).Appendix A The tight-binding model
@@ A.3 TB+U
A.4 Goodness of fitting the Hubbard-
To quantify our model’s power to reproduce the reference results we calculated the following expression for layer number and stacking configurations:
where are the magnetic moments from [47] and [9] and are the magnetic moments from our work. This value can be seen in Fig. 9.
@@ Appendix C Simultaneous uniaxial strain and compression
In [47] and [9] the effect of different geometries was not studied. In accordance with this we fixed the lattice parameters. With these settings the TB+U model could not reproduce the absence of magnetism in 3, 4 and 7 layer ABA graphene. Calculation incorporating both uniaxial strain and interlayer compression (Fig. 10) show that for a large set of possible lattice parameters the magnetic moments are negligibly small or zero for the ABA 3 layer graphene. The same effect occurs in 5 layer ABA but for a much smaller set of lattice parameters. Meanwhile the magnetism is robust for both ABC multilayers. Therefore an explanation for the absence of magnetism in the 3,4 and 7 layer ABA systems in [9] is the selected values of the lattice parameters. Finally, we emphasize a limitation of the present approach: our TB+ model employs a single on-site (purely local) effective Hubbard parameter within a mean-field treatment. This neglects nonlocal Coulomb interaction terms (e.g., intersite interactions) and the screening/environment dependence of the effective interaction. Such beyond-local effects are expected to be most relevant when the system is close to the critical point and the magnetic moments are extremely small, as in ABA trilayer/tetralayer/heptalayer.